Startup Night 5 - 20140903 UT

Observers: JHill, BRothberg (LBTO), OKuhn, TGolota, AConrad (Tucson)
Telescope Operator: GBechetti (LBTO)
AO Support: JCGuerra
Telescope Support: MWagner, JLittle (LBTO)
Software Support: MDeLPena
Instrument Support: ESolheid (LBTO)
Engineering Support: DAshby

Summary

Photometric skies, with seeing varying from 0.45 to 1.2 arcsec during the night.

Non-Sidereal Guding with MODS1

This evening we conducted a successful test of using guide stars with MODS1 to track 3 asteroids moving relative to the fixed stars. To guide for the asteroid, GCS has to move the guide probe with a reflex motion in the opposite direction. We did 3 asteroids: 49 Pales moving at 7 arcsec/hour, 80 Sappho moving at 27 arcsec/hour and 1943 Anteros moving at 114 arcsec/hour. We also took a MODS1 spectrum of 1943 Anteros.

Thanks to Geno, Barry, Olga, Al and Taras who assisted. While this was a great success, it wasn't quite perfect. The asteroid has a sawtooth jitter with an amplitude of about 2 arcsec. I'm about 80% sure I know the explanation for this jitter.

AGW1 wake-up

Summary: The AGW1 wake-up has gone fine tonight, so from the AGW1 perspective we are ready to have LUCI1 mounted tomorrow.

Details: We fought with AGW1 and oacserver for about 45 minutes with the usual sort of RPC errors and Network errors. Once we got that straightened out, everything was fine. The AGw1 gave 0.5 arcsec images right off the bat. I've adjusted the probe to the rotator center, and am now collecting data for Doug to measure the pupil position.

WRS testing with LBC

We got 2 hours of WRS testing. We did get it to converge to decent images on Red, but there are still some issues related to the rotation of astigmatism and trefoil.

Preparations

New version of IIF - to update collimation zernike LUT on preset. v8.26a

New version of PSF - to use the last good temperature for collimation. v6.7b

Running GCS v2.17.4 for MODS1 NS guiding

Details

01:54 UT (and earlier) - Testing LBC - bias, test images, darks.

01:55 UT - Getting ready to open

02:10 UT - Waking up MODS, Enclosure open

02:16 UT - Taking Sieve mask snaps

02:25 UT - collimating

Asteroid 49 Pales

Hijacked NS ACTIVE preset to asteroid 49 Pales with guidestars from Pales_20140903_0300.catalog. rate=7 arcsec/hour

02:37 UT - test exposures for MODs w/non-sidereal - seeing 1.3" via Guider

02:43 UT - taking series of 10 x10sec exposures in g and r with MODS

Details for sending the NS preset

See NonSiderealGuiding

Asteroid 80 Sappho

Hijacked NS ACTIVE preset to asteroid 80 Sappho with guidestars from Sappho_20140903_0300.catalog. rate=27 arcsec/hour

02:52 UT - moving to a faster asteroid - stopped blue at 0010, red at 0011 (0011 is bad - telescope is moving)

02:55 UT - starting series of 10exposures for Sappho

03:01 UT - stopped red at 16 and blue at 15

03:02 UT - changing to shorter exposure time (3sec blue and red)

03:03 UT - resent presets and non-sidereal, 0018 (red) and 0017 (blue) are streaked (unusable)

03:15 UT - Starting another set of 10 exposures (blue and red, 0027 & 0028, respectively)

03:25 UT - left preset cancelled - but still in set of 10 exposures

03:27 UT - Finished 2nd set on Sappho (don't use blue 0036 and red 0037 - telescope moved while integrating)

Asteroid 1943 Anteros

03:28 UT - Hijacked NS ACTIVE preset to 1943 Anteros, rate = 114 arcsec/hour

03:33 UT - on target - taking test 30sec exposure (0037 blue, 0038 red) - saturated asteroid -switching to 10sec

03:37 UT - 0038 blue, 0039 red 10sec each - red saturated, blue counts at 30k - going to 5sec each

03:39 UT - 0039 blue, 0040 red - 5sec - looks good

03:41 UT - 15exposures, 5sec each blue (0040) and red (0041) started - seeing 1.3" Guider, 1.1" DIMM - did not do IMCS lock again - this may have caused jumps before

03:58 UT - starting second set of 15 exposures blue & red (g and r)

04:18 UT - guiding stopped

04:21 UT - restart guiding - will attempt spectra of asteroid 1943 Anteros with 1.2" slit dual

04:31 UT - 1.2" dual grating spectra of 1943 Anteros - 600sec in blue and red

04:44 UT - faint continuum seen in red side - taking thru-slit image

04:49 UT - did not see asteroid in thru-slit - taking field image

04:55 UT - aligned on wrong target - reacquired correct object (asteroid)

04:56 UT - taking another 600 sec dual 1.2" spectrum of asteroid to test sidereal tracking long term

04:57 UT - lost guiding - stopped integration

05:00 UT - sending new preset with 0500 ephemeris

05:12 UT - acquired on asteroid - trying another 600sec dual (0080 red, 0072 blue)

05:26 UT - taking thru slit image to check (red 0081)- slipped from Y=1682 to Y=1677

05:28 UT - taking field images (red 0082, 0083, 0084, 0085) to check it matches thru-slit and Y position of spectra

05:34 UT - field images show slight jitter, but asteroid appears to have remained in the slit for the most part

05:43 UT - Switching to test AGW1

AGW1 wake-up at LFBG

Authorize failed because SX M3 wasn't initialized.

Lots of problems getting AGW1 and GCS to communicate.

06:20 Cold boot of agw-control0

06:25 Call Taras - still another restart of oacserver does the trick.

06:32 ACQUIRE preset succeeds IE=-54 CA=-32

06:36 ACTIVE preset to BS9183 gs=0 - collimation succeeds at 0.45-0.5 arcsec FWHM on the guider.

06:43 ACTIVE preset to BS9183 gs=5 (mag=15) - off-axis image was immediately 0.5 arcsec.

AGW1 hotspot to rotator center alignment

http://wiki.lbto.org/bin/view/Commissioning/BGTechProcedures

data directory is /home/lbto/data/20140903_agw1

06:47 ACTIVE preset to BS9183 gs=0

stop guiding

rotator to 0

ptincrement 0 IE 5

setxy -u 1 -x 0 -y 612.5

rotator to 360

readGuideCam left -e 80000 left_guider000002.fits center is 113,66 (binned by 2) ccdsec1 is 8:260, so add 8,15 --> 121,81, multiply by 2 --> 242,162

[telescope@agw-control0 ~]$ getxy -u 1 -o 16
Guide probe position
	x: -1.000000
	y: 1001660.000000

getHotspot left Hotspot is (x,y): 255.00, 230.00

Is the hotspot correct? No, spots are at the bottom of the field stops. So we need to adjust hotspot......while guiding continues.

setHotspot left -x 255 -y 200 -- thats too high in wfsc 0315 (more negative y moves the star image up on wfsc)

07:16 setHotspot left -x 255 -y 215 -- just a hair low in wfsc 0321

07:19 setHotspot left -x 265 -y 215 (more positive x moves the star image left on wfsc)

07:22 setHotspot left -x 260 -y 214 which looks OK in wfsc 0328 THIS IS THE NEW HOTSPOT.

Want to increase star image by 18 pixels in X = 0.9 arcsec = 0.5 mm, so decrease ROT homeoffset by 30
Want to increase star image by 50 pixels in Y = 2.5 arcsec = 1.5 mm, so decrease RAD ooffset by 1500

as agwuser

vi oacontrol.conf

stop GCS

rdwrconfig -u 1
call failed: RPC: Timed out
rdwrconfig: rdwrconfig failed (AGW1: RPC Communication Error)
(probably because Geno restarted GCS too quickly)

07:43 stop and restart oacserver, AGW1, GCS

Editted new hotspot into LUCI_L.cfg

07:47 ACTIVE preset to BS9183 gs=0

stop guiding

rotator to 360

setxy -u 1 -x 0 -y 612.5

rotator to 0

readGuideCam left -e 80000 left_guider000003.fits star is a point at 121,89 --> 129,104 --> 258,208

ptincrement 0 IE 5

rotator to 360

readGuideCam left -e 80000 left_guider000004.fits center 121,87 so add 8,15 --> 129,102, multiply by 2 --> 258,204

We are calling this good.

Transform Data on StoneA1

08:05 ACTIVE preset to StoneA1 gs=0 - no star found

Adjusted pointing, still no star found for StoneA1. ???

iif_register, /TCS

Transform Data on StoneA2

08:13 ACTIVE preset to StoneA2 gs=0

08:14 transform_collect for StoneA2_new

silly John, you can't take transform data without LUCI......................

Field Aberration Data on BS9107

IDL does not have the Hotspot or Pupil Positions updated, but this data set will let Doug measure the pupil positions.

08:18 field_collect on BS9107 - acquisition missed

08:21 ACQUIRE preset to adjust pointing

08:23 field_collect on BS9107 BS9107_new field list

Seeing is about 0.9 arcsec on guider, but wfs rms variations are around 100, so local seeing must be small. Seeing is degrading as this data set goes on, then improving again near the end.

#| GStar | wfsc # | 201493 BS9107_new_field wfsc GCS
0 368-370
23 371-373
0 374
28 375-377
0 378
36 379-381
0 382-383
33 384-386
0 387-388
18 389-391
0 392-393
39 394-396
0 397-398
24 399-401
0 402-403
42 404-406
0 407
20 408-410
0 411-412
50 413-415
0 416-417
39 418-420
0 421-423
43 424-426

Sending a couple presets to get off the Y-axis pupils

09:23 BS9107 gs=43 PA=-20 121.8,123

09:25 BS9107 gs=43 PA=68 -119,5,126.1

WRS Tests with LBCs

09:29 To zenith for reconfiguration to LBC on both sides.

09:51 LBC preset to WT10_342

WRS notes are on the WRS wiki page.

LBC Standards (or are they sky flats?)

Sky is photometric, but getting bright. Blue is not well collimated.

12:04 Galadi1_4p_Vgr -- mainly to see if the rotator is OK. Sky is getting bright. Only Blue images.

12:09 Galadi1-UBVRI (stopped halfway through)

Close

12:16 Close for impending dawn

12:17 1 pair of bias frames.



-- %BUBBLESIG{BarryRothberg - 2014-09-03}%
Topic revision: r8 - 22 Oct 2014, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback